A wheel 1.0 m in radius rotates with an angular acceleration of 4.0rad/s2


1) If the wheel’s initial angular velocity is 2.0 rad/s, its angular velocity after 10 s is


2)The angle in radians it rotates in the 10-s interval is


3)The tangential speed of the point on the rim of the wheel at the end of the 10-s interval is


4)The tangential acceleration of the point on the rim of the wheel at the end of the 10-s interval is


1
Expert's answer
2021-03-17T16:52:49-0400

1) If the wheel’s initial angular velocity is 2.0 rad/s, its angular velocity after 10 s is


"\omega_f=\omega_i+\alpha t=2+4·10=42\text{ rad/s}."

2) The angle in radians it rotates in the 10-s interval is


"\phi=\omega_it+\frac{\alpha t^2}{2}=220\text{ rad}."

3) The tangential speed of the point on the rim of the wheel at the end of the 10-s interval is


"v=\omega_fR=42·1=42\text{ m/s}."

4) The tangential acceleration of the point on the rim of the wheel at the end of the 10-s interval is


"a=\alpha R=4·1=4\text{ m/s}^2."